OpenStudy (anonymous):

Lara tossed a fair coin 3 times. What is the probability of getting heads in the first two trials? 1 over 8 2 over 8 3 over 8 4 over 8

OpenStudy (anonymous):

@paki

OpenStudy (anonymous):

@mayankdevnani @sleepyjess

OpenStudy (anonymous):

I think D

OpenStudy (mayankdevnani):

@77777jeannie77777 can you made events ? i mean HHH, TTT etc,

OpenStudy (anonymous):

I'm not really sure how to do that :(

OpenStudy (mayankdevnani):

make events

OpenStudy (mayankdevnani):

1) HHH 2) HHT 3) HTH 4) THH 5) HTT 6) THT 7) TTH 8) TTT \[\large \bf total~no.~of~trials=2^3=8\]

OpenStudy (mayankdevnani):

now, you have to choose that trials or events in which first two trials you get head :- 1) HHH 2) HHT only 2

OpenStudy (mayankdevnani):

so \[\large \bf P(E)=\frac{no.of~trials}{Total~no.of~outcomes}=\frac{2}{8}=\frac{1}{4}\]

OpenStudy (anonymous):

oh so then its 1/2 probability?

OpenStudy (mayankdevnani):

hope you understand. @77777jeannie77777

OpenStudy (mayankdevnani):

no @77777jeannie77777 you have to calculate :- `probability of getting heads in the first two trials`

OpenStudy (mayankdevnani):

understood?? @77777jeannie77777

OpenStudy (anonymous):

hmm..... wait a sec plz lemme see

OpenStudy (mayankdevnani):

:)

OpenStudy (anonymous):

OH so then its umm 1/4 probability??

OpenStudy (mayankdevnani):

correct !

OpenStudy (anonymous):

If its that then the answer is B, right??

OpenStudy (mayankdevnani):

right !

OpenStudy (anonymous):

@mayankdevnani ??

OpenStudy (mayankdevnani):

you are right !

OpenStudy (mayankdevnani):

@77777jeannie77777 ?? lol

OpenStudy (anonymous):

lol sorry thx :):)